If the pressure in a cylinder of carbon dioxide at 34°C increases from 1.02 atm to 1.34 atm what is the new temperature of the gas?
1
Expert's answer
2020-06-22T09:16:59-0400

Solution.

"T_1=34^oC=307K;"

"P_1=1.02atm;"

"P_2=1.34atm;"

"V=const;"

"\dfrac{P_1}{T_1}=\dfrac{P_2}{T_2}\implies" "T_2=\dfrac{T_1P_2}{P_1};"


"T_2=\dfrac{307K\sdot1.34atm}{1.02atm}=403K=130^oC;"


Answer: "T_2=130^oC."